# Ultrabroadband Resonant Frequency Doubling on a Chip

### v1

@@ Abstract
Microresonators are powerful tools to enhance the efficiency of second-order nonlinear optical processes, such as second-harmonic generation, which can coherently bridge octave-spaced spectral bands. However, dispersion constraints such as phase-matching and doubly resonant conditions have so far limited demonstrations to narrowband operation. In this work, we overcome these limitations showing ultrabroadband resonant frequency doubling in a novel integrated device, wherein the resonant enhancement of pump and second harmonic are individually addressed in two distinct and linearly uncoupled microring resonators, each adjusted to target the respective spectral band. The two microresonators are designed and tuned independently, yet share a common interaction region that grants nonlinear coupling over a quasi-phase-matching bandwidth exceeding 200 nm, enabled by the inscription of a photoinduced grating. The system allows to not only conveniently disentangle the design parameters of the two microresonators but also to reconfigure the doubly resonant condition electrically, and the phase-matching condition optically. We demonstrate milliwatt-level addressable second-harmonic generation over the entire telecom band and then configure the device to generate and upconvert a Kerr frequency comb with bandwidth exceeding 100 nm and upconverted power up to 10 mW.
@@
While microring resonators are capable of highly efficient frequency doubling, their practical usage is hindered by limitations in bandwidth and poor wavelength tunability. Indeed, broadband SHG in a ring resonator must satisfy three requirements simultaneously: (i) phase-matching, (ii) doubly resonant condition for pump and SH, and (iii) free spectral range (FSR) matching. Perfect phase-matching is usually prevented by material and modal dispersion, but it can be obtained through intermodal phase-matching [19] or through quasi-phase-matching (QPM) — the latter either exploiting the natural birefrengence of the medium (-QPM) [11] or the electric-field poling of a ferroelectric waveguide [9]. Doubly resonant high-Q microrings are also typically unachievable deterministically, even by tailored designs, due to fabrication tolerances, for the resonance position must be controlled with a precision comparable with the linewdith. A common workaround is thermal tuning [24], where the slightly different thermo-optic effect on pump and SH resonances can compensate for small deviations from a nearly doubly resonant condition statistically obtained among many fabricated samples. However, this method is not scalable, does not allow for the precise control of the operation wavelength, and becomes unpractical in systems with large free spectral range (FSR). If a doubly resonant condition is difficult to achieve at a given pump wavelength, broadband SHG involving multiple resonances is even more challenging. Indeed, the FSRs at pump and SH frequencies are in general different: therefore, high-Q resonators only enable doubly resonant SHG over a small number of resonances at the time, restricting applications such as frequency comb upconversion to a limited number of spectral lines [25, 26] and, more generally, setting a trade-off between bandwidth and CE [27, 28, 29]. In principle, this can be mitigated by adjusting the waveguide cross-section to match the group velocity [21], but at the cost of dispersion engineering, which is already required for controlling the comb properties or generating soliton states.
@@
In this work, we address all of these issues, inherent to the single-ring approach, by demonstrating a device where the pump and SH fields are associated with two separate resonators that yet share a common path where nonlinear processes can occur. This change of paradigm allows one to adjust the optical properties of the system in different spectral regions independently and achieve efficient nonlinear interaction without dispersion engineering [30, 31, 32, 33, 34, 35]. First, a photoinduced nonlinearity is realized in the Si3N4 coupling region where pump and SH coexist, whose periodicity automatically satisfies the QPM condition over a broad bandwidth (all-optical poling process, AOP [20]). Second, the families of azimuthal modes of the two resonators can be independently tuned through electric actuators to make sure that a doubly resonant condition is always satisfied. Third, the microresonator lengths are tailored to grant FSR matching at the design level. Finally, efficient in-coupling and extraction of light at the pump and SH wavelength are obtained by optimizing the couplers to the two resonators independently. We investigate the device operating principle by two-photon microscopy, showing that the AOP process occurs in the region where the circulating pump and SH field interact, proving the linearly uncoupled nature of the two resonators. We then showcase the potential of this approach by reconfiguring the device to achieve SHG with milliwatt-level generated power over the C and L telecom bands, limited only by the bandwidth of our amplifiers. Finally, we demonstrate the broadband operation of the device by using it to generate a modulation instability comb that we frequency-double within a pump bandwidth of nearly ().
@@ Results
A conceptual schematic of the reconfigurable SHG source is shown in Fig. 1a. The device is composed of two “racetrack” resonators designed to operate respectively in the pump (“south”) and SH (“north”) spectral bands. Both resonators are critically coupled to their bus waveguide for the fundamental transverse-electric mode (TE00). A Mach-Zehnder interferometer (MZI) structure is used to introduce an interaction region between the two resonators by using two directional couplers designed to operate near 100:0 splitting ratio in the pump band and near 0:100 in the SH band [35]. Finally, a thermo-optic phase shifter (“heater”) is placed on each resonator and on the lower arm of the MZI, allowing, respectively, to tune the set of resonant modes at pump/SH and to compensate for any residual linear coupling between the two racetracks.
@@
A typical fabricated device is shown in Figs. 1c-d. We use a lensed fiber to couple light to the chip and electrical probes to actuate the heaters (see Methods). Typical transmission spectra in the pump and SH bands are shown in Fig. 1e. By acting on the south and north resonator heaters we are able to observe a rigid and independent shift of the respective resonance spectra (the north heater case is shown inf Fig. 1e). Similarly, acting on the MZI heater enables us to dynamically vary the pump resonances visibility and linewidth, as a result of power leakage to the north resonator, strongly overcoupled at the FH (see Supplementary Note 1). Once these values are optimized, from best fit of the Lorentzian lines recorded we estimate a loaded Q factor of around at the pump and at the SH.
@@
To investigate SHG, we pump the south resonator with an amplified tunable laser and configure the device with the procedure described in the Methods. The occurrence of the AOP process is marked by a sudden increase of the SHG power from the north resonator. Visible and infrared imaging of the device in operation (Fig. 2a) provides straightforward evidence of the linearly uncoupled nature of the process. Light scattering from the north and south resonators during the SHG process, when acquired with either a visible or an infrared camera, suggests that the SH and pump fields are circulating only in their respective resonators. A typical spectrum of the output light is also shown in Fig. 2b, confirming the SHG process and the absence of spurious spectral components. By repeating this procedure, we reproduce a similar result among several devices with different FSR and coupling conditions, systematically recording a generated SH power of up to .
We then focus our attention on a single device, characterized by a nominal FSR of about , and investigate its suitability for efficient SHG across multiple resonances. To this end, we tune the pump wavelength to each of the resonances within the C and L telecom bands (1530-1620 nm) and repeat the optimization procedure, each time reconfiguring the QPM grating by AOP. Fig. 2c shows the recorded SHG spectra, all corresponding to an output power in the milliwatt level (average: ), proving the capability of the integrated nonlinear device to operate across a wideband range when electrically reconfigured. Note that this demonstration is limited by the availability of amplifiers, while the nonlinear device is expected to be capable of covering an even broader spectrum, limited only by the bandwidth of the resonators’ point couplers and of the MZI, estimated at about ().
@@
The experiments described so far entail circulating pump and SH intensities sufficiently high to trigger the AOP process. However, it is known that the grating, once inscribed, can subsist for very long time, exceeding several months at room temperature [36]. We therefore investigate the nonlinear response of the system in a regime of low pump power, insufficient to trigger the erasure or reconfiguration of the grating, after previous poling a selected resonance. To establish the limits of such “cold cavity” regime, we perform a scaling analysis of the generated SH power as a function of the pump power in the bus waveguide, as shown in Fig. 3b. For each value of the pump power, the wavelength is slowly scanned through the resonance from the blue to the red side, and the peak SHG power is recorded. The resulting trend shows a quadratic increase of the generated SH for pump powers below approximately . When the pump power is further increased, a deviation from such trend is recorded: we attribute this to the reconfiguration of the grating. When further increasing the pump power, the generated SH restores the quadratic trend, reentering the combined AOP/SHG process (“hot cavity” regime). The cold- and hot-cavity regimes are therefore desirable for efficient SHG, while the intermediate regime of reconfiguration should be avoided.
@@ Discussion
[…]
The shaded area highlights silicon nitride photonic platforms. The bandwidth is estimated at 10 dB; power and CE are intended at CW regime unless otherwise stated. n.a.: data not available.
@@
The device developed represents the first demonstration of addressable doubly resonant SHG through external tuning, overcoming several fundamental limitations inherent in the single-resonator approach. Our strategy can be generally set in the context of coupled resonator systems (or “photonic molecules”) which have been recently employed to improve the performance of other nonlinear processes, such as comb generation [42], self-injection locking [43], and quantum light generation [44, 45]. However, our approach differs fundamentally from the above due to the deliberate choice of eliminating the linear coupling, with the advantages of (i) strongly suppressing the parasitic loss due to overcoupling of north resonator’s modes at the pump (ii) avoiding the formation of supermodes, associated with a lower field enhancement (equivalently, larger mode volume) and (iii) avoiding (or controlling) alterations to the dispersion profile associated with mode anti-crossings.
This strategy, originally introduced in the context of spontaneous parametric sources [30, 31, 32, 33, 34] and here combined with AOP, enables electrical control over the doubly resonant condition within a pump bandwidth exceeding (), only limited in this demonstration by the range of our amplifiers. The system can systematically generate milliwatt-level powers, with CE as high as 40%/W in CW regime. These values are already useful for several practical applications and can likely be improved through further reduction of the propagation loss and by achieving a better rejection of the residual linear and intermodal coupling at the pump, further mitigating the need for high pump power.
@@ Acknowledgements
This work was funded by the European Research Council grant PISSARRO (ERC-2017-CoG 771647) and by the Swiss National Science Foundation (grant 214889). L.Z. and M.L. acknowledge support from PNRR MUR project PE0000023-NQSTI.
Competing interests The authors declare no competing interests.
Data availability
The data and code that support the plots within this paper and other findings of this study are available from the corresponding authors upon reasonable request.

### v2

# Ultrabroadband Milliwatt-Level Resonant Frequency Doubling on a Chip

@@ Abstract
Microresonators are powerful tools to enhance the efficiency of second-order nonlinear optical processes, such as second-harmonic generation, which can coherently bridge octave-spaced spectral bands. However, dispersion constraints such as phase-matching and doubly resonant conditions have so far limited demonstrations to narrowband operation. In this work, we overcome these limitations showing ultrabroadband resonant frequency doubling in a novel integrated device, wherein the resonant enhancement of pump and second harmonic are individually addressed in two distinct and linearly uncoupled microring resonators, each adjusted to target the respective spectral band. The two microresonators are designed and tuned independently, yet share a common interaction region that grants nonlinear coupling over a quasi-phase-matching bandwidth exceeding 200 nm, enabled by the inscription of a photoinduced grating. The system allows to not only conveniently disentangle the design parameters of the two microresonators but also to reconfigure the doubly resonant condition electrically, and the phase-matching condition optically. We demonstrate milliwatt-level addressable second-harmonic generation over the entire telecom band and then configure the device to internally generate and upconvert a Kerr frequency comb with bandwidth exceeding 100 nm and upconverted power up to 10 mW.
Introduction
@@
While microring resonators are capable of highly efficient frequency doubling, their practical usage is hindered by limitations in bandwidth and poor wavelength tunability. Indeed, broadband SHG in a ring resonator must satisfy three requirements simultaneously: (i) phase-matching, (ii) doubly resonant condition for pump and SH, and (iii) free spectral range (FSR) matching. Perfect phase-matching is usually prevented by material and modal dispersion, but it can be obtained through intermodal phase-matching [19] or through quasi-phase-matching (QPM) — the latter either exploiting the natural birefringence of the medium (-QPM) [11] or the electric-field poling of a ferroelectric waveguide [9]. Doubly resonant high-Q microrings are also typically unachievable deterministically, even by tailored designs, due to fabrication tolerances, for the resonance position must be controlled with a precision comparable with the linewidth. A common workaround is thermal tuning [24], where the slightly different thermo-optic effect on pump and SH resonances can compensate for small deviations from a nearly doubly resonant condition statistically obtained among many fabricated samples. However, this method is not scalable, does not allow for the precise control of the operation wavelength, and becomes unpractical in systems with large FSR. If a doubly resonant condition is difficult to achieve at a given pump wavelength, broadband SHG involving multiple resonances is even more challenging. Indeed, the FSRs at pump and SH frequencies are in general different: therefore, high-Q resonators only enable doubly resonant SHG over a small number of resonances at the time, restricting applications such as frequency comb upconversion to a limited number of spectral lines [25, 26] and, more generally, setting a trade-off between bandwidth and CE [27, 28, 29]. In principle, this can be mitigated by adjusting the waveguide cross-section to match the group velocity [21], but at the cost of dispersion engineering, which is already required for controlling the comb properties or generating soliton states.
@@
In this work, we address all of these issues, inherent to the single-ring approach, by demonstrating a device where the pump and SH fields are associated with two separate resonators that yet share a common path where nonlinear processes can occur. This change of paradigm allows one to adjust the optical properties of the system in different spectral regions independently and achieve efficient nonlinear interaction without dispersion engineering [30, 31, 32, 33, 34, 35]. First, a photoinduced nonlinearity is realized in the Si3N4 coupling region where pump and SH coexist, whose periodicity automatically satisfies the QPM condition over a broad bandwidth (all-optical poling process, AOP [20]). Second, the families of azimuthal modes of the two resonators can be independently tuned through electric actuators to make sure that a doubly resonant condition is always satisfied. Third, the microresonator lengths are tailored to grant FSR matching at the design level. Finally, efficient in-coupling and extraction of light at the pump and SH wavelength are obtained by optimizing the couplers to the two resonators independently. We investigate the device operating principle by two-photon microscopy, showing that the AOP process occurs in the region where the circulating pump and SH field interact, proving the linearly uncoupled nature of the two resonators. We then showcase the potential of this approach by reconfiguring the device to achieve SHG with milliwatt-level generated power over the C and L telecom bands, limited only by the bandwidth of our amplifiers. Finally, we demonstrate the broadband operation of the device by using it to generate a modulation instability comb that we frequency-double intracavity, within the same structure, over a pump bandwidth of nearly ().
@@ Results
A conceptual schematic of the reconfigurable SHG source is shown in Fig. 1a. The device is composed of two “racetrack” resonators designed to operate respectively in the pump (“south”) and SH (“north”) spectral bands. Both resonators are nearly critically coupled to their bus waveguide for the fundamental transverse-electric mode (TE00). A Mach-Zehnder interferometer (MZI) structure is used to introduce an interaction region between the two resonators by using two directional couplers designed to operate near 100:0 splitting ratio in the pump band and near 0:100 in the SH band [35]. Finally, a thermo-optic phase shifter (“heater”) is placed on each resonator and on the lower arm of the MZI, allowing, respectively, to tune the set of resonant modes at pump/SH and to compensate for any residual linear coupling between the two racetracks.
@@
A typical fabricated device is shown in Figs. 1c-d. We use a lensed fiber to couple light to the chip and electrical probes to actuate the heaters (see Methods). Typical transmission spectra in the pump and SH bands are shown in Fig. 1e. By acting on the south and north resonator heaters we are able to observe a rigid and independent shift of the respective resonance spectra (the north heater case is shown in Fig. 1e). Similarly, acting on the MZI heater enables us to dynamically vary the pump resonances visibility and linewidth, as a result of power leakage to the north resonator, strongly overcoupled in the pump band (see Supplementary Note 1). Once these values are optimized, from best fit of the Lorentzian lines recorded we estimate a loaded Q factor of around at the pump and at the SH.
@@
To investigate SHG, we pump the south resonator with an amplified tunable laser and configure the device with the procedure described in the Methods. The occurrence of the AOP process is marked by a sudden increase of the SHG power from the north resonator. Visible and infrared imaging of the device in operation (Fig. 2a) provides straightforward evidence of the linearly uncoupled nature of the process. Light scattering from the north and south resonators during the SHG process, when acquired with either a visible or an infrared camera, suggests that the SH and pump fields are circulating only in their respective resonators. A typical spectrum of the output light is also shown in Fig. 2b, confirming the SHG process and the absence of spurious spectral components. By repeating this procedure, we reproduce a similar result among several devices with different FSR and coupling conditions, systematically recording a generated SH power of up to (with a pump power of ).
We then focus our attention on a single device, characterized by a nominal FSR of about , and investigate its suitability for efficient SHG across multiple resonances. To this end, we tune the pump wavelength to each of the resonances within the C and L telecom bands (1530-1620 nm) and repeat the optimization procedure, each time reconfiguring the QPM grating by AOP. Fig. 2c shows the recorded SHG spectra, all corresponding to an output power in the milliwatt level (average: 1.4 dBm, standard deviation: 2.4 dB, see Supplementary Figure 6 for more details), proving the capability of the integrated nonlinear device to operate across a wideband range when electrically reconfigured. Note that this demonstration is limited by the availability of amplifiers, while the nonlinear device is expected to be capable of covering an even broader spectrum, limited only by the bandwidth of the resonators’ point couplers and of the MZI, estimated at about ().
@@
The experiments described so far entail circulating pump and SH intensities sufficiently high to trigger the AOP process. However, it is known that the grating, once inscribed, can subsist for very long time, exceeding several months at room temperature [36]. We therefore investigate the nonlinear response of the system in a regime of low pump power, insufficient to trigger the erasure or reconfiguration of the grating, after previous poling a selected resonance. To establish the limits of such “cold cavity” regime, we perform a scaling analysis of the generated SH power as a function of the pump power in the bus waveguide, as shown in Fig. 3b. For each value of the pump power, the wavelength is slowly scanned through the resonance from the blue to the red side, and the peak SHG power is recorded. The resulting trend shows a quadratic increase of the generated SH for pump powers below approximately . When the pump power is further increased, a deviation from such trend is recorded: we attribute this to the reconfiguration of the grating [37]. When further increasing the pump power, the generated SH restores the quadratic trend, reentering the combined AOP/SHG process (“hot cavity” regime). The cold- and hot-cavity regimes are therefore desirable for efficient SHG, while the intermediate regime of reconfiguration should be avoided. At higher power, we expect the quadratic trend to eventually saturate owing to the increased nonlinear photoconductivity [20, 18].
@@ Discussion
| Si3N4 waveguide [15] | 0.05 | 4.5 | 0.16 | QPM (all-optical poling) |
| Si3N4 waveguide [21] | 0.15 | >25 | n.a. | QPM (all-optical poling) |
| Si3N4 waveguide [17] | 0.003 | >100 | 0.6c | QPM (all-optical poling) |
| Si3N4 microring [25] | 0.16 | 10b | 1 | Intermodal |
| Si3N4 microring [19] | 2,500 | 0.06a,d | 2.2 | Intermodal |
| Si3N4 microring [20] | 47.6 | 0.6a,d | 12.5 | QPM (all-optical poling) |
| Si3N4 microring [24] | 2 | 2.3 | 2 | QPM (all-optical poling) |
| Si3N4 microresonators system (this work) | 40 | >100b | >10 | QPM (all-optical poling) |
The bandwidth is estimated at 10 dB; power and CE are intended at CW regime unless otherwise stated. n.a.: data not available.
@@
The device developed represents the first demonstration of addressable doubly resonant SHG through external tuning, overcoming several fundamental limitations inherent in the single-resonator approach. Our strategy can be generally set in the context of coupled resonator systems (or “photonic molecules”) which have been recently employed to improve the performance of other nonlinear processes, such as comb generation [42], self-injection locking [43], and quantum light generation [44, 45]. However, our approach differs fundamentally from the above due to the deliberate choice of eliminating the linear coupling, with the advantages of (i) strongly suppressing the parasitic loss due to overcoupling of north resonator’s modes at the pump (ii) avoiding the formation of supermodes, associated with a lower field enhancement (equivalently, larger mode volume) and (iii) avoiding (or controlling) alterations to the dispersion profile associated with mode anti-crossings. It’s also worth noticing that our paradigm shares similarities with hybrid racetrack geometries recently employed for the demonstration of optical parametric oscillators in TFLN [46, 47], where a non-resonant mode at short wavelength is linearly uncoupled, but nonlinearly coupled with the resonant modes of a microresonator at longer wavelengths.
Our strategy, originally introduced in the context of spontaneous parametric sources [30, 31, 32, 33, 34] and here combined with AOP, enables electrical control over the doubly resonant condition within a pump bandwidth exceeding (), only limited in this demonstration by the range of our amplifiers. The system can systematically generate milliwatt-level powers, with CE as high as 40%/W in CW regime. These values are already useful for several practical applications and can likely be improved through further reduction of the propagation loss and by achieving a better rejection of the residual linear and intermodal coupling at the pump, further mitigating the need for high pump power. Note that the CE could be further improved by increasing the resonator quality factors without compromising the phase-matching bandwidth, thanks to the flexible engineering of resonance positions. Alternative designs for linear uncoupling (see, e.g., Ref. [35]) could also help mitigate the current efficiency penalty, albeit at the expense of reduced QPM bandwidth (see Methods).
@@
A further perspective advantage of our approach is the possibility of continuously tuning the doubly-resonant SHG. By synchronizing a pump laser scan with the tuning of the south and north heaters, one could potentially achieve a “gap-free” tuning of the generated SH field, which could be particularly beneficial to address specific SH wavelengths, such as those associated with atomic transitions. Essential requirements to this end are (i) the possibility to continuously tune the pump and SH resonances across at least one FSR (here around , realistically achievable with standard heater modules) and (ii) a seamless transition between different pairs of resonances involved in the AOP-SHG process, that should be slow enough to allow the reconfiguration of the grating.
@@ Methods
Data availability
The data and code that support the plots within this paper and other findings of this study are available from the corresponding authors upon request.
@@ Acknowledgements
M.C., J.Z., and C.-S.B. acknowledge funding by the European Research Council grant PISSARRO (ERC-2017-CoG 771647) and by the Swiss National Science Foundation (grant 214889). M.L. acknowledges PNRR MUR project “National Quantum Science and Technology Institute” – NQSTI (Grant No. PE0000023).
Author contributions
M.C. led the project, designed the devices, performed the measurements, and carried out the analysis and physical interpretation of the results. L.Z. and M.L. contributed to the device design and theoretical modeling. J.Z. performed preliminary characterization of the experimental devices. M.C., M.L., and C.-S.B. conceived the original idea. All authors participated in the analysis and discussion of the results. M.C. wrote the manuscript with assistance from C.-S.B. and contributions from all authors. C.-S.B. supervised the project.
Competing interests
The authors declare no competing interests.